OpenStudy (anonymous):

I need help with this question as well.. I got the correct slope and all but idk why I'm getting a different answer? If someone could check my work please that would be great :)

OpenStudy (anonymous):

Here is my work! (Attached a picture)

OpenStudy (irishboy123):



OpenStudy (irishboy123):

use y = mx + c you have m and you have a point , ie an (x,y) on the curve, so go that way

OpenStudy (irishboy123):

\(\sqrt{3} = 3 \times \frac{2}{\sqrt{3}} + c\)

OpenStudy (irishboy123):

and repeat for the normal after the -1/m step

OpenStudy (anonymous):

Thank you!!! That was very helpful:)